Margarita operates a sole proprietorship that earns $100,000 of qualified business income after deducting salaries of $300,000. The sole proprietorship is not a specified service business. She files a single tax return for 2019. Assume her taxable income before the QBI deduction is $175,000. Margarita's QBI deduction for 2019 is:

a.$20,000.

b.$-0-.

c.$60,000.

d.$80,000.

e.$35,000.
Answer #1

a) $ 20,000

Qbi deductionis lesser of

i) 20% QBI income = $ 100,000 * 20% = $ 20,000

II) 20% Of taxable income less capital gains = $ 175,000 * 20%= $ 35,000

As taxable income is more than $ 157,500, QBI deduction is limited to greater of :

i) $ 300,000 * 50% = $ 150,000

II) 25% * 300,000 + 2.5% * 0% = $ 75,000